OpenStudy (thecatman):

Solve the equation t over 3 equals negative 3. t equals negative nine t equals negative one t equals one t equals nine

OpenStudy (timaashorty):

this is the equation: \[\frac{ t }{ 3 } = -3\] So to find What t equals you need to do the opposite. Multiply 3 by -3. So \[3 x -3 = -3.\] you're answer will then be the first one. >> t equals negative nine

OpenStudy (timaashorty):

**** Oops I ment: \[3 x -3 = -9\] not -3.

OpenStudy (thecatman):

ok thanks

OpenStudy (timaashorty):

your welcome.